### v1

@@ 5.1 Implications and limits of the argument
In the general example presented at the beginning of Sec. 4, we mentioned non-tested hypotheses. In this perspective, one could think that the EP is instead a well tested hypothesis. However, the argument considers not the EP itself, but the fundamentality of the EP, which is not decisively tested. In fact, especially at quantum level, the EP is tested but not decisively tested. The situation would instead radically change if a fundamental theory, requiring EP at any level, were formulated.
On the other hand, one could include, in the argument, all theories predicting the violation of the EP, in order to show that also these theories disagree with the hypothesis $FEP$. However, these theories are not really dinamically equivalent to GR as in the case of TEGR and STEGR. Even if some of them would recover the GR phenomenology, at today status of art, they cannot be considered the final theory of gravity. They may have good empirical and observational evidences in different regimes and scales, as in the case of MOND, but they give not the same predictions $P_{1},...,P_{n}$. Therefore, the same epistemic considerations would not be valid.
@@
The first two problems $(i)$ and $(ii)$ can be addressed thanks to the Palatini formalism, where the metric $g_{\mu\nu}$ and the connection $\Gamma^{\rho}_{\mu\nu}$ are independent, with $g_{\mu\nu}$ determining the causal structure and $\Gamma^{\rho}_{\mu\nu}$ the free fall, i.e. the geodesic structure. In this way, the connection becomes the true fundamental dynamical variable and the observable of the theory. Following the words by Einstein, the metric is ”dethroned” and becomes an ”ancillary variable” (see Ref. [6] and references therein).
The tension with new physics predictions $(iii)$, with quantum mechanics $(iv)$ and the unjustified a priori coincidence between $m_{I}$ and $m_{G}$ $(vii)$ would be resolved, since the EP could be considered as an emergent feature, recoverable but not postulated. So the only requirements would be the compatibility of theories with experimental constraints.
Then, TEGR and STEGR allow gravity to be reformulated as a gauge theory $(vi)$, and their extensions $f(T)$ and $f(Q)$ could be studied in order to search for a geometrical solution to DE and CDM problems $(v)$.
Finally, without imposing the EP, we are not privileging the spacetime structure over torsion or non-metricity a priori (viii). The geometric under-determination would remain, but the correct spacetime structure would no longer be determined by a postulate.
@@
Second, it could be objected that the number of true predictions which would be considered sufficient in order to trust the further predictions of the theories is somewhat arbitrary. Are $four$ good predictions sufficient? Actually, in this case, there are five good predictions, since the $EMEP$ is a true non-trivial and very significant prediction of the theories. One could point out that predictions $S$ and $C$ are actually consequences of prediction $FE$, since both the solutions and the cosmological applications depend on the field equations. This is true; maybe they are not independent predictions, but the interesting thing is again that with the extensions $f(T)$ and $f(Q)$, the equivalence is not anymore valid [9], so it is important to explore any possible different feature in solutions and applications [48, 49].
@@ 5.2 Experimental perspectives
As, mentioned, physicists are not yet satisfied with the current precision of the EP tests, as we can see from the numerous experiments which are developed by many different research groups [21]. These tests are complex and require many years of work and experimental efforts, not to mention the proposed space missions. All this efforts are developed in order to increase the accuracy of the EP tests, demonstrating the unsatisfactory situation, especially at quantum level.
Our argument does not mean that EP is false or that it has to be necessarily violated at some level, but surely it encourages experimentalists in the search for a possible violation to discriminate among concurring theories of gravity. If EP is not fundamental, it could be an emergent property. Therefore, it could be the case that it is violated at some level, for example at quantum level. The incoming experiments of free falling with quantum tests could be the straightforward approach to probe the above statement (see for instance [62, 21, 63]). Clearly these quantum tests are WEP tests, and $\neg$¬SEP does not imply $\neg$¬WEP, but $\neg$¬WEP does imply $\neg$¬SEP, i.e. if we detect a violation of the WEP, this would falsify also the SEP. This is relevant because the finest experiments on EP we have at the moment are conceived for the weak formulation. Clearly, the free fall of a wave packet is something different with respect to the free fall of a classical test particle. This conceptual aspect needs further and deep investigations.
Currently, the highest accuracy on the Eötvös parameter, which quantify the violation of the WEP, has been reached by the MICROSCOPE space mission with a free-fall experiment performed with macroscopic classical masses. In 2017, they reached $10^{-14}$ [64] and, in 2022, $10^{-15}$ [65]. Other future space missions have been proposed, such as the Galileo Galilei (GG) [66] and the Satellite Test of the Equivalence Principle (STEP) [67], with the goal of achieving $10^{-17}$ and $10^{-18}$, respectively. The atomic experiments have compared the free fall of different isotopes or atomic species such as 85Rb and 87Rb, 39K and 87Rb, the bosonic 88Sr and the fermionic 87Sr and also atoms in different spin orientations. In an experiment in Stanford [68], a precision of $10^{-12}$ was reached (see [21, 63] for a review). Similarly to the classical counterparts, the ultimate performance of atomic sensors for WEP tests can be reached in space, where tests with a precision of $10^{-15}\div 10^{-17}$ were proposed by the STE-QUEST (Space–Time Explorer and QUantum Equivalence Space Test) mission [69, 70]. Experiments exploiting entangled atomic states aim to push the sensitivity beyond the so called Standard Quantum Limit [71]. Fig. 4 shows the limits set by the WEP tests performed with different methods, from 1960 until today and the future prospects.
As for the SEP tests, the achieved accuracy limit is $\eta_{SEP}\approx 10^{-5}$ [21].
@@
If a violation of the EP were to be detected, it would constitute a falsification of GR at that level, while TEGR and STEGR would remain viable as theories. Specifically, such a finding would suggest that the EP is not fundamental ($FEP$) but rather an emergent property ($EMEP$) arising from a possible gauge choice, consistent with TEGR and STEGR. In this case, it would mean that the EP is recoverable only at limited levels, with constraints highlighted by experiments. This is analogous to Einstein’s justification of the EP on empirical grounds. In neither case do we have fundamental reasons to assume the EP. However, GR explicitly assumes it as fundamental, with all the consequences we have discussed, whereas TEGR and STEGR do not.
@@
However, it is important to emphasize that this argument encourages testing the EP but does not strictly predict its violation. In fact, even if the EP is not fundamental, it is still possible that no violation will be found. In such a case, increasing the accuracy of traditional EP tests may be insufficient to discriminate among the three theories.
@@
This consideration suggests the need to explore and design other types of experiments that could highlight the difference in essence between gravity and inertia, investigate the possible distinct dynamics of the metric and the connection, or determine in other way whether the EP is fundamental or emergent. Consequently, these epistemic considerations point to the importance of developing experimental approaches beyond traditional EP tests, which might potentially differentiate among the three equivalent theories. This represents a subject for future research.
@@
Then note that the continue corroboration of the EP at any level poses also other conceptual problems. If after significant progress, no violation will be found, at what level of accuracy would we consider it satisfied? $10^{-18}$? $10^{-20}$? $10^{-50}$? What level of precision could be deemed sufficient to corroborate a metric theory? This is a perfect example of inductive risk, as one could always find an experiment which violates the EP, even if we will reach a precision of $10^{-50}$. Is there an accuracy level where reasonable doubt would be mitigated? Maybe finding a more fundamental theory that explains why the EP should be an $FEP$ could help. However, justifying it experimentally seems difficult, as it is possible that its validity might turn out to be merely a contingent fact without any underlying fundamental reason. Philosophers of physics surely would enjoy the debate.
@@ 6 Conclusions
We discussed that EP is a non-trivial, theory-laden assumption in the framework of Equivalent Gravities. In fact, assuming EP at the foundation of the theory, we are intrinsically stating that gravitation is given by the curvature of spacetime, rather than by torsion or non-metricity. This fact gives rise to many problematic foundational consequences: $i)$ the coincidence between the causal and the geodesic structure; $ii)$ the fact that metric tensor is the fundamental variable of the theory instead of the connection; $iii)$ the contrast with several new physics predictions; $iv)$ the conceptual difficulties at quantum level; $v)$ the relation with CDM and DE problems; $vi)$ the conceptual difference of gravity with respect to other gauge theories; and, finally, $vii)$ the unjustified coincidence between gravitational and inertial mass. All these difficulties (or shortcomings) are direct or indirect consequences of the assumption that gravitation and inertia are of the same essence, in the words of Einstein.
@@
Then, on a closer inspection, we argued that there is a crucial hidden difference in relation with EP. The significant fact is that, in both TEGR and STEGR, EP can be recovered but without the necessity to postulate it at the foundation of the theory. If physicists would have found the EP as a necessary fundamental principle also for TEGR and STEGR, this would have been considered as a meta-empirical evidence in favor of what we called the FEP, that is the Equivalence Principle as Fundamental. But, as we have seen, this is not the case. Therefore, given the equivalence among GR, TEGR and STEGR in non-trivial multiple predictions, and given the fact that EP is not necessary for TEGR and STEGR, our confidence in the fundamentality of the EP decreases.
If the argument is correct, this could represent a difference in the empirical content between the three theories, because if the EP is not a fundamental feature of reality (FEP), it is emergent (EMEP). And if it is emergent, it is possible that, at some level, it is not valid.
@@
These epistemic considerations encourage physicists to further enhance the accuracy of EP tests, especially at the quantum level, and to develop new experimental schemes to investigate potential differences in the empirical content of theories arising from their distinct relationships with the EP, such as discriminating between dynamics of metric and connection.
Finally, the approach of Sec. 4 could be used also for other investigations. First of all, it should be used to evaluate other predictions in order to search for other possible relevant observables. For example, principles as the Local Lorentz Invariance, the Local Position Invariance, that together with EP constitute the Schiff conjecture, should be investigated also in TEGR and STEGR. Then, for instance, there could be also other equivalent formulations of gravity outside of the Geometric Trinity. Finding other dynamically equivalent theories would help in further constraining the configuration space of the theory. Having already found two of such theories, nothing precludes the fact that other equivalent representations of gravity could exist. Beside EP, theories of gravity could be compared also considering the number of degrees of freedom related to observables. This will be the argument of further studies.
@@
Acknowledgements C.M. and S.C. acknowledge the support of INFN sez. di Napoli, iniziative specifiche QGSKY and MOONLIGHT2. S.C. thanks the Gruppo Nazionale di Fisica Matematica of Istituto Nazionale di Alta Matematica for the support. This paper is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). The Authors thank Elena Castellani, Carmen Ferrara, and Flaminia Giacomini for the useful discussions and feedbacks.

### v2

@@ 5.1 Implications and limits of the argument
In the general example presented at the beginning of Sec. 4, we mentioned non-tested hypotheses. In this perspective, one could think that the SEP is instead a well tested hypothesis. However, the argument does not address the SEP itself, but rather the fundamentality of SEP, which has not been decisively tested — as other authors have suggested, questioning the universality of EP rather than its fundamentality, using similar arguments based on teleparallel gravity [64, ch. 9]. In fact, particularly at quantum level, the WEP, which is at the foundation of SEP, is tested but not decisively tested [72]. The situation would instead radically change if a fundamental theory, requiring some form of EP at any level, were formulated.
On the other hand, one could include, in the argument, all theories predicting the violation of the SEP, in order to show that also these theories disagree with the hypothesis $FEP$. However, these theories are not really dynamically equivalent to GR as in the case of TEGR and STEGR. Even if some of them would recover the GR phenomenology, given the state of the art today, they cannot be considered the final theory of gravity. They may have good empirical and observational evidences in different regimes and scales, as in the case of MOND, but they give not the same predictions $P_{1},...,P_{n}$. Therefore, the same epistemic considerations would not be valid.
@@
The first two problems $(i)$ and $(ii)$ can be addressed thanks to the Palatini formalism, where the metric $g_{\mu\nu}$ and the connection $\Gamma^{\rho}_{\mu\nu}$ are independent, with $g_{\mu\nu}$ determining the causal structure and $\Gamma^{\rho}_{\mu\nu}$ the free fall, i.e. the geodesic structure. In this way, the connection becomes the true fundamental dynamical variable and the observable of the theory. Following the words by Einstein, the metric is “dethroned” and becomes an “ancillary variable” (see Ref. [6] and references therein).
The tension with new physics predictions $(iii)$, with quantum mechanics $(iv)$ and the unjustified a priori coincidence between $m_{I}$ and $m_{G}$ $(vii)$ would be resolved, since the SEP, along with the WEP and the EEP, could be considered an emergent feature, recoverable but not postulated. So the only requirements would be the compatibility of theories with experimental constraints. Then, TEGR and STEGR allow gravity to be reformulated as a gauge theory $(vi)$, and their extensions $f(T)$ and $f(Q)$ could be studied in order to search for a geometrical solution to DE and CDM problems $(v)$.
Finally, without imposing the SEP, we are not privileging the spacetime structure over torsion or non-metricity a priori (viii). The geometric under-determination would remain, but the correct spacetime structure would no longer be determined by a postulate.
@@
Second, it could be objected that the number of true predictions which would be considered sufficient in order to trust the further predictions of the theories is somewhat arbitrary. Are fourfour good predictions sufficient? Actually, in this case, there are five good predictions, since the $EMEP$ is a true non-trivial and very significant prediction of the theories. One could point out that predictions $S$ and $C$ are actually consequences of prediction $FE$, since both the solutions and the cosmological applications depend on the field equations. This is true; maybe they are not independent predictions, but the interesting thing is again that with the extensions $f(T)$ and $f(Q)$, the equivalence is not anymore valid [9], so it is important to explore any possible different feature in solutions and applications [53, 54].
@@ 5.2 Experimental perspectives
As mentioned, physicists are not yet satisfied with the current precision of the tests of the different formulations of the EP, as we can see from the numerous experiments which are developed by many different research groups [26]. These tests are complex and require many years of work and experimental efforts, not to mention the proposed space missions. All this efforts are developed in order to increase the accuracy of the EP tests, demonstrating the unsatisfactory situation, especially at quantum level.
[…]
Currently, the highest accuracy on the Eötvös parameter, which quantify the violation of the WEP, has been reached by the MICROSCOPE space mission with a free-fall experiment performed with macroscopic classical masses. In 2017, they reached $10^{-14}$ [75] and, in 2022, $10^{-15}$ [76]. Other future space missions have been proposed, such as the Galileo Galilei (GG) [77] and the Satellite Test of the Equivalence Principle (STEP) [78], with the goal of achieving $10^{-17}$ and $10^{-18}$, respectively. The atomic experiments have compared the free fall of different isotopes or atomic species such as 85Rb and 87Rb, 39K and 87Rb, the bosonic 88Sr and the fermionic 87Sr and also atoms in different spin orientations. In an experiment in Stanford [79], a precision of $10^{-12}$ was reached (see [26, 74] for a review). Similarly to the classical counterparts, the ultimate performance of atomic sensors for WEP tests can be reached in space, where tests with a precision of $10^{-15}\div 10^{-17}$ were proposed by the STE-QUEST (Space–Time Explorer and QUantum Equivalence Space Test) mission [80, 81]. Experiments exploiting entangled atomic states aim to push the sensitivity beyond the so called Standard Quantum Limit [82].
Finally, although the technology is not yet mature for high-precision measurements, there has been progress in developing experiments to test the WEP with antimatter [26, 83, 84]. This line of research is particularly interesting, as a WEP test using a matter-wave interferometry with antihydrogen or positronium would probe both the quantum and antimatter regimes, domains in which the validity of the WEP remains uncertain.
Fig. 4 shows the limits set by the WEP tests performed with different methods, from 1960 until today and the future prospects. For the SEP tests in the sense of Will [4, p. 76], the achieved accuracy limit is $\eta_{SEP}\approx 10^{-5}$ [26].
@@
If a violation of some form of the EP were to be detected, it would constitute a falsification of GR at that level, while TEGR and STEGR would remain viable as theories. This point is controversial. Some authors argue that, since the three nodes of the geometric trinity are dynamically equivalent, no empirical differences between them should exist [18]. According to this view, a violation of the SEP would falsify not only GR, but also TEGR and STEGR at that level. However, GR postulates that the SEP is fundamental and holds at every point in spacetime, whereas in TEGR and STEGR it can be recovered through possible gauge choices. For them, in principle there may exist points where the SEP does not hold. If experiments were to reveal that the SEP fails at some point, the gauge choices that restore the SEP at those regions would constitute surplus structure, unnecessary for representational purposes [18]. This, however, would not amount to a falsification of TEGR or STEGR. On the contrary, such a result could be seen as supporting them, as some proponents of TEGR have argued [64, ch. 9]. Specifically, such a finding would suggest that the SEP is not fundamental ($FEP$) but rather an emergent property ($EMEP$) arising from a possible gauge choice, consistent with TEGR and STEGR. In this case, it would mean that the SEP is recoverable only at limited levels, with constraints highlighted by experiments. This is analogous to Einstein’s justification of the EEP on empirical grounds. In neither case do we have fundamental reasons to assume the given EP. However, GR explicitly assumes it as fundamental, with all the consequences we have discussed, whereas TEGR and STEGR do not. Moreover, recall that, as Wolf &\& Read argue, the empirical content of a theory is not fully determined by its dynamics [19], and theories sharing the same action may exhibit differences in their empirical content.
Another interesting possibility, suggested by the Read and Teh taxonomy [24], concerns the fact that a falsification of WEP would not imply a falsification of SEP2,¬EEP, although its recovery should still be constrained by experimental data. In this scenario as well, TEGR would be a more fundamental theory than GR. Both these two scenario would be in line with the suggestion by Weatherall that both TEGR and STEGR have more surplus structure than GR [18].
@@
However, it is important to emphasize that this argument encourages testing the SEP but does not strictly predict its violation. In fact, even if the SEP is not fundamental, it is still possible that no violation will be found. In such a case, increasing the accuracy of traditional EP tests may be insufficient to discriminate among the three theories.
@@
This consideration suggests the need to explore and design other types of experiments that could highlight the difference in essence between gravity and inertia, investigate the possible distinct dynamics of the metric and the connection, or determine in other way whether the SEP is fundamental or emergent. Consequently, these epistemic considerations point to the importance of developing experimental approaches beyond traditional EP tests, which might potentially differentiate among the three equivalent theories. This represents a subject for future research.
@@
Then note that the continue corroboration of the SEP at any level poses also other conceptual problems. If after significant progress, no violation will be found, at what level of accuracy would we consider it satisfied? $10^{-18}$? $10^{-20}$? $10^{-50}$? What level of precision could be deemed sufficient to corroborate a metric theory? This is a perfect example of inductive risk, as one could always find an experiment which violates the EP, even if we will reach a precision of $10^{-50}$. Is there an accuracy level where reasonable doubt would be mitigated? Maybe finding a more fundamental theory that explains why the SEP should be an $FEP$ could help. However, justifying it experimentally seems difficult, as it is possible that its validity might turn out to be merely a contingent fact without any underlying fundamental reason. Philosophers of physics surely would enjoy the debate.
@@ 6 Conclusions
We discussed that SEP is a non-trivial, theory-laden assumption in the framework of Equivalent Gravities. In fact, assuming SEP at the foundation of the theory, we are intrinsically stating that gravitation is given by the curvature of spacetime, rather than by torsion or non-metricity. This assumption also leads to a number of other problematic foundational consequences: $i)$ the coincidence between the causal and the geodesic structure; ii)ii) the fact that metric tensor is the fundamental variable of the theory instead of the connection; $iii)$ the contrast with several new physics predictions; $iv)$ the conceptual difficulties at quantum level; $v)$ the relation with CDM and DE problems; $vi)$ the conceptual difference of gravity with respect to other gauge theories; and, finally, $vii)$ the unjustified coincidence between gravitational and inertial mass. All these difficulties (or shortcomings) are direct or indirect consequences of the imposition of SEP, which is based on the assumption that gravitation and inertia are of the same essence (as expressed by EEP).
@@
Then, on a closer inspection, we argued that there is a crucial hidden difference in relation with SEP. The significant fact is that, in both TEGR and STEGR, SEP can be recovered but without the necessity to postulate it at the foundation of the theory. If physicists would have found the SEP as a necessary fundamental principle also for TEGR and STEGR, this would have been considered as a meta-empirical evidence in favor of what we called the FEP, that is the Equivalence Principle as Fundamental. But, as we have seen, this is not the case. Therefore, given the equivalence among GR, TEGR and STEGR in non-trivial multiple predictions, and given the fact that SEP is not necessary for TEGR and STEGR, our confidence in the fundamentality of the SEP decreases.
If the argument is correct, this structural divergence could lead to differences in the empirical content between the three theories, because if the SEP is not a fundamental feature of reality (FEP), it is emergent (EMEP). And if it is emergent, it is possible that, at some level, it is not valid.
@@
These epistemic considerations encourage physicists to further enhance the accuracy of EP tests, especially at the quantum level and unexplored regimes, and to develop new experimental schemes to investigate potential differences in the empirical content of theories arising from their distinct relationships with the EP, such as discriminating between dynamics of metric and connection.
Finally, the approach of Sec. 4 could be used also for other investigations. First of all, it should be used to evaluate other predictions in order to search for other possible relevant observables. For example, principles as the Local Lorentz Invariance, the Local Position Invariance, the Schiff conjecture, should be investigated also in TEGR and STEGR. Then, for instance, there could be also other equivalent formulations of gravity outside of the Geometric Trinity. Finding other dynamically equivalent theories would help in further constraining the configuration space of the theory. Having already found two of such theories, nothing precludes the fact that other equivalent representations of gravity could exist. Beside EP, theories of gravity could be compared also considering the number of degrees of freedom related to observables. This will be the argument of further studies.
@@
Acknowledgements The Authors thank Elena Castellani, Carmen Ferrara, and Flaminia Giacomini for the useful discussions and feedbacks. C.M. and S.C. acknowledge the support of INFN sez. di Napoli, iniziative specifiche QGSKY and MOONLIGHT2. S.C. thanks the Gruppo Nazionale di Fisica Matematica of Istituto Nazionale di Alta Matematica for the support. This paper is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).